# What happens when generative AI models train recursively on each others’ generated outputs?

### v2

@@ Abstract
The internet is full of AI-generated content while also serving as a common source of training data for generative AI (genAI) models. This duality raises the possibility that future genAI models may be trained on other models’ generated outputs. Prior work has studied consequences of models training on their own generated outputs, but limited work has considered what happens if models ingest content produced by other models. Given society’s increasing dependence on genAI tools, understanding downstream effects of such data-mediated model interactions is critical. To this end, we provide empirical evidence for how data-mediated interactions might unfold in practice, develop a theoretical model for this interactive training process, and show experimentally possible long-term results of such interactions. We find that data-mediated interactions can benefit models by exposing them to novel concepts perhaps missed in original training data, but also can homogenize their performance on shared tasks.
@@ 1 Introduction
Since the release of ChatGPT in 2022, generative AI (genAI) models have exploded in popularity. Now capable of generating highly realistic text, images, and videos, these models have been widely adopted for various use cases, from creative idea generation [4, 28] to healthcare support [52] to national security settings [33, 30]. Given the significant uptick in genAI use across numerous industries, it is clear that this technology is here to stay. Consequently, interrogating potential ways genAI models could evolve—in positive or harmful ways—is critical.
@@
With few exceptions, today’s large-scale genAI models are trained on massive datasets sourced from the internet. Widely-accepted scaling laws for model performance say that training on more data aids learning[40], and the internet provides a rich, cheap, and ever-evolving source of training data. Although whitepapers for more recent genAI models withhold details about training set composition—potentially due to ongoing litigation about copyright concerns—evidence from earlier whitepapers indicates that scraped data was used to train models like Llama, Gemini, Phi, the GPT series, Claude, and others [27, 7, 58, 38, 3, 14, 5].
@@
Beyond privacy and copyright concerns, training on scraped data could have other downsides. Prior work has noted genAI models trained recursively on their own generated outputs “collapse,” becoming unable to generate meaningful content [56, 34, 46, 11]. This scenario is feasible, since AI-generated content abounds online [57] and could be part of future training datasets. However, subsequent work has proposed ways to mitigate collapse via reuse of non-AI-generated data in subsequent training iterations [23, 41, 24, 29]. Model collapse remains an activate research area [55].
@@
Yet, prior work studying the dynamics of model collapse has overlooked another reality: the internet teems with content from many genAI models. Today’s most popular models have millions of users [53, 32, 8], who leverage generative AI tools to create online content like web pages and social media posts [1]. Recent work showed that up to 40% of content on popular sites like Quora is now AI-generated [57]. Given the increasing availability of these models for a variety of public-facing uses [2, 4, 28], AI-generated content from many different models will continue to proliferate.
@@
The standard practice of training on scraped internet data and the increasing prevalence of AI-generated content online suggest the strong possibility that future generative AI models will be trained on other models’ outputs. Yet, this aspect of model training has received considerably less attention in prior work. Given the widespread adoption of generative AI models in critical settings like healthcare and national security, this phenomenon and its downstream effects ought to be vetted to ensure models remain helpful and trustworthy.
@@
Contributions. To address this need, this work theoretically derives and experimentally investigates the long-term evolutionary behavior of models trained on each other’s data. Specifically, we
@@
Both our theoretical analysis and experiments show that when training with a mixture of real and synthetic data, the implicit interaction between heterogeneous models and datasets can have both positive and negative impacts. At a high level, the reasons for this can be anticipated from well known concepts in statistical learning theory: recursive training on the same data is bad (e.g., overfitting and model “collapse”) but training on novel data, even if synthetic, can boost performance (e.g., transfer learning). Our experimental results provide concrete evidence that these phenomena can occur simultaneously. Our theory shows that the full range of observed behaviors can be modeled accurately via a linear dynamical system.
@@ 2 Related Work
Model collapse is a recently observed phenomenon in large-scale generative text and image models. It referred—in it’s earliest form—to the phenomenon of models performing much worse when trained on their own generated outputs [56, 49, 64, 11, 29, 34, 24, 46]. Theoretical and empirical results from these work show that models, if trained on generated outputs from their prior versions, slowly degrade in performance as generations progress. One way this manifests is in models forgetting the tails of their original (real) training data, since generated content tends not to contain rare content from the original training data. Training repeatedly on truncated, synthetic data leads the model to forget the richness of its original distribution, resulting in degraded behavior (at best) and total failure (at worst).
Mitigating model collapse. Despite the dire predictions of these works, subsequent work has proposed a simple mitigation strategy: instead of discarding all prior (human-generated) training data, retain some fraction of this while augmenting it with generated data. Numerous works have observed that this choice to augment instead of discard the original training dataset results in a bounded error in future models, avoiding collapse [41, 31, 45]. Although most of these results were discovered on small models, recent work claims that the observed bound in error $\pi^{2}/6$ exists for all models [23]. Further work [55] summarizes current research on collapse and calls for greater clarity and precision in discussing this phenomenon.
Transfer learning and other model interactions. Significant prior work has studied the phenomenon of transfer learning, in which information learned by one model is passed to another, often by reusing the trained weights of a “teacher” model to initialize a “student” model [71]. Some prior work has further considered the use of synthetic data in transfer learning [61, 42, 15]. Our work is distinct from transfer learning due to its focus on unintentional data-mediated interactions between models. Furthermore, limited work has specifically examined long-term effects of models intentionally training on each other’s data. One work [69] considers the setting where a generative model is trained on data generated by other models, but does not consider long-term effects of such interactions among multiple models. Another recent work [37] studies interacting Large Language Model agents (e.g. LLMs with extra capabilities like web browsing) through the lens of Bayesian social learning and microeconomics, but does not focus specifically on data-mediated interactions between models.
@@ 3 How Today’s Large-Scale Generative AI Models Are Trained
Step 1: Pretraining. Following well-established scaling laws linking model performance and dataset size [40], today’s large-scale generative AI models are trained on massive datasets often scraped from the internet. Early versions of GPT, Llama, and PaLM all report being trained on scraped datasets like Common Crawl, ArXiv, Github, Wikipedia, and/or Stack Exchange [62, 17, 16]. Table 1 summarizes publicly documented use of scraped internet data in model training.
@@
Chinchilla [36]
GLaM [26]
GPT [50]
GPT-2 [51]
GPT-3 [16]
LaMDA [60]
Llama 1 [62]
PaLM [17]
Phi 2 [6]
Table 1: Examples of training data listed for prominent language models. = use explicitly stated; = significant overlap expected (e.g. GLaM [26] and PaLM [17] are trained on Microsoft’s internal re-creation of WebText [16]). We only include models for which training data sources are explicitly stated. See Table 3 in Appendix for information on other prominent models.
@@
Another striking fact emerges from the categorization of training data in Table 1: large-scale model training datasets overlap. For example, GPT, Jamba, Llama, PaLM, and Phi are all trained on subsets of CommonCrawl [21], while GPT, Llama, and PaLM are all trained on Wikipedia and Books datasets. Several other models have other points of training data overlap. This suggests that:
@@
This overlap in training datasets has not been meaningfully explored in the literature and may have important implications for how models evolve.
Step 2: Fine-tuning. Variously called fine-tuning or alignment, this phase leverages proprietary methods or data to tweak model behaviors in ways model providers believe are helpful. For example, the LLama fine-tuning phase [27] involves many rounds of reinforcement learning with human feedback (RLHF) to stamp out model negative behaviors, while Phi [5] was fine-tuned on proprietary synthetic data tailored to cover “gaps” in original corpus related to mathematical reasoning.
Step 3: Model updates. A key assumption of prior literature on model collapse is that models are updated. Updating generally involves training a new model using similar architecture initialized with weights from the prior step, techniques, and datasets to prior versions of the model. This matches reality—the literature documents numerous “families” of models that directly descend from one another (c.f. LLama 1, 2, and 3 [62, 63, 27]; GPT 1, 2, 3, 4 [50, 16, 51, 7]; Phi 1.5, 2, 3 [43, 6, 5]). Prior works on model collapse have proposed several ways training data from prior model generations could be re-used (or not) during model updates:
  • The replace scenario, first proposed in [56] assumes that a model trainer completely replaces their training data at each update, using only generated data output by the prior version of the model. As recent work has pointed out, this edge case is interesting but impractical [55].
  • The accumulate scenario [31] assumes model trainers augment their original training data with additional data that may contain AI-generated content. Dataset grows linearly at each update.
  • Finally, the accumulate and subsample scenario [41] echoes the accumulate scenario but assumes a bounded dataset size. It subsamples a fixed subset from the original and accumulated data at each model update. This scenario acknowledges the real-world compute limitations model trainers face.
@@
[62]: ArXiv, Books, Common Crawl,
@@
[27]: "A variety of data sources containing
@@
GPT [50]: BooksCorpus [51]: WebText
[16]: CommonCrawl, WebText2,
@@
[43]: The Stack, Stack Overflow,
@@
[6]: The Stack, Stack Overflow,
@@
[5]: "publicly available web data…
@@
We believe that the accumulate-and-subsample paradigm best reflects reality, so we will leverage it in our work. We support this opinion with evidence from three well-documented model families: Llama, GPT, and Phi. Table 2 records the training data used in publicly disclosed generations of these models. As the table shows, trainers re-use some prior training data for model updates, supplementing this with additional content derived from the web. Whitepapers for models published after 2023 do not contain specific training data information, perhaps because of pending litigation over copyright concerns in training data use, e.g. [48], making it difficult to trace data re-use after this point. However, the language in these papers suggests collection of additional online data at each update.
@@
However, most prior work on model collapse still overlooks a fundamental reality in model updates: future models trained on internet-sourced content will be trained on outputs from other generative AI models, not merely their own. Already, the internet is filled with generated content from various models [57]. As model trainers collect new data to facilitate model updates, internet-sourced data will inevitably contain content from other generative models.
@@ 4 Formalizing An Iterative, Interactive Model Training Pipeline
Section 3 provides empirical evidence for two realities overlooked by prior studies: internet-scraped training datasets used for initial training may have substantial overlap, and models may be updated using each others’ generated outputs. Both factors could significantly influence model evolution, and it is imperative to understand how they influence models long-term. To study the effect of these two factors on model evolution, we propose a general workflow in which multiple entities regularly update their models using a mix of private, public, and generated data. Based on the evidence of §3, we consider three basic types of training/update data:
@@
  • $D_{t}=\{D_{t1},D_{t1},\ldots D_{tk}\}$: Public data used for updates at time $t$ by multiple entities (synthetic data). $D_{tk}$ is data generated by the $k^{th}$ entity based on model $\theta_{t-1,k}$.
@@
  • $\alpha$, $0\leq\alpha\leq 1$ is the fraction of new data introduced at generation $t$, relative to the amount of initial data reused (following the “accumulate and subsample” paradigm of prior work [41]).
@@
    Each entity $k=1,\dots,K$ updates it model parameters via $\hat{\theta}_{t,k}=\Phi_{t,k}(\hat{\theta}_{t-1,k},D_{*},\tilde{D}_{k},D_{t})$ Here, $\Phi_{0,k}$ is a training algorithm that depends on the previous model parameter $\hat{\theta}_{k,t-1}$ as well as the data. As before, training may employ subsampling, weighting, and randomization.
@@ 5 Theory
We theoretically analyze the behavior of the interactive workflow. Similar to prior work [31, 41, 23, 24] we focus on the linear regression models where each data point consists of a feature-response pair $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$. By the universality results of [23], the analysis of this settings also applies to generalized linear models satisfying appropriate asymptotic normality assumptions.
@@
Throughout our analysis we assume that all features are deterministic. We represent dataset $\tilde{D}_{k}$ with $\tilde{n}_{k}\times d$ matrix $\tilde{X}_{k}=[\tilde{x}_{k1},\dots,\tilde{x}_{k\tilde{n}_{k}}]^{\top}$ and $\tilde{n}_{k}\times 1$ vector $\tilde{y}_{k}=[\tilde{y}_{k1},\dots,y_{k\tilde{n}_{k}}]^{\top}$, and use the same convention for the public data $(X_{*},y_{*})$ and the generated data $(X_{tk},y_{tk})$. Data across different entities are then combined into “lifted” representations, which are denoted using boldface:
@@
Classical results in matrix analysis [35] imply that if the spectral radius of $\bm{Q}$ is strictly less than one, then these these matrices converge to well-defined limits $\bm{M}$ and $\bm{C}$ satisfying
@@
The following result provides a sufficient condition for convergence in terms of the triple $(\tilde{\bm{S}},S_{*},\bm{S})$. In particular, if $\bm{S}$ is proportional to $\tilde{\bm{S}}$ then the condition is satisfied for all $10\leq\alpha<$ 1 and $0\leq\beta\leq 1$. Note that the boundary case $\alpha=1$ corresponds to the recursive training setting in [56] where the variance increases linearly with the number of generations, and thus convergence does not occur.
@@
Suppose that $\bm{S}\propto\lambda\tilde{\bm{S}}+(1-\lambda){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\mathrm{I}_{K}\otimes S_{*})}$ for some $0<\lambda\leq 1$. Then, the spectral radius of $\bm{Q}$ is strictly less than one for all $10\leq\alpha<$ 1 and $0<\beta\leq\lambda$.
@@
Figure 2 compares the asymptotic MSE for training workflow (obtained from Theorem 3) with the MSE of an idealized setting where each entity has access to the entire collection of real data (both private and public). Each curve represents the relative efficiency, i.e., the ratio of optimal MSE to entity-specific workflow MSE, with values close to one indicating near optimality. These results, which corresponds to a single realizations of i.i.d. standard Gaussian features in $d=15$ dimensions, demonstrate that small to moderate values of $\alpha$ can provide global improvements over the initialization ($\alpha=0$) with higher diversity between entities (larger $\beta$) resulting in more dramatic improvements.
@@
Figure 2: Relative efficiency across $\alpha$ values for a $K=4$ model system at varying $\beta$. Each curve is the ratio between the MSE of the minimum variance unbiased estimator based on all the real data (both private and public) and the asymptotic MSE obtained from Theorem 3. We use dimension $15$ and rank $5$.
6 Experiments with Large Language Models
To understand how our theoretical predictions bear out in practice, we perform experiments with $K$ different LLMs. In all cases, the high-level task is to answer natural language questions with high fluency and accuracy. The “private” datasets associated with each of the models correspond to specific subdomains such as science based questions [39, 18] or grade school level math problems [20]. We follow the training workflow outlined in Section 4. Results for $K=3$ are presented Figure 4, and additional results are provided in the supplementary material.

### v3

# What happens when generative AI models train recursively on each others’ outputs?

@@ Abstract
The internet serves as a common source of training data for generative AI (genAI) models but is increasingly populated with AI-generated content. This duality raises the possibility that future genAI models may be trained on other models’ generated outputs. Prior work has studied consequences of models training on their own generated outputs, but limited work has considered what happens if models ingest content produced by other models. Given society’s increasing dependence on genAI tools, understanding such data-mediated model interactions is critical. This work provides empirical evidence for how data-mediated interactions might unfold in practice, develops a theoretical model for this interactive training process, and experimentally validates the theory. We find that data-mediated interactions can benefit models by exposing them to novel concepts perhaps missed in original training data, but also can homogenize their performance on shared tasks.
@@ 1 Introduction
Since the release of ChatGPT in 2022, generative AI (genAI) models have exploded in popularity. Now capable of generating highly realistic text, images, and videos, these models have been widely adopted for various use cases, from creative idea generation (Notion, 2024; Ellis, 2024) to healthcare support (Reddy, 2024) to national security settings (Harding, 2024; GAO, 2024). Given the significant uptick in genAI use across numerous industries, this technology is clearly here to stay. Consequently, interrogating potential ways genAI models could evolve—in positive or harmful ways—is critical.
@@
With few exceptions, today’s large-scale genAI models are trained on massive datasets sourced from the internet. Widely-accepted scaling laws for model performance say that training on more data aids learning (Kaplan et al., 2020), and the internet provides a rich, cheap, and ever-evolving source of training data. Although whitepapers for more recent genAI models withhold details about training set composition—potentially due to ongoing litigation about copyright concerns—evidence from earlier whitepapers indicates that scraped data was used to train models like Llama, Gemini, Phi, the GPT series, Claude, and others (Dubey et al., 2024; Achiam et al., 2023; Team et al., 2024a; Jiang et al., 2023; coh, 2024; Anthropic, 2023; Abdin et al., 2024).
@@
Beyond privacy and copyright concerns, training on scraped data could have other downsides. Prior work has noted that genAI models trained recursively on their own generated outputs “collapse,” becoming unable to generate meaningful content (Shumailov et al., 2024; Hataya et al., 2023; Martínez et al., 2023; Alemohammad et al., 2024). This scenario is feasible, since AI-generated content abounds online (Sun et al., 2025) and could be part of future training datasets. However, subsequent work has proposed ways to mitigate collapse via reuse of non-AI-generated data in subsequent training iterations (Dey & Donoho, 2024; Kazdan et al., 2025; Dohmatob et al., 2025; Feng et al., 2024). Model collapse remains an activate research area (Schaeffer et al., 2025).
@@
Yet, prior work studying the dynamics of model collapse has overlooked another reality: the internet teems with content from many genAI models. Today’s most popular models have millions of users (Reuters, 2024; Handa et al., 2025; AI, 2024a), who leverage generative AI tools to create online content like web pages and social media posts (gen, 2022). Recent work showed that up to 40% of content on popular sites like Quora is now AI-generated (Sun et al., 2025). Given the increasing availability of these models for a variety of public-facing uses (app, 2024; Notion, 2024; Ellis, 2024), AI-generated content from many different models will continue to proliferate.
@@
The standard practice of training on scraped internet data and the increasing prevalence of AI-generated content online suggest the strong possibility that future generative AI models will be trained on other models’ outputs. Yet, this aspect of model training has received relatively little attention. Given the widespread adoption of generative AI models in critical settings like healthcare and national security, this phenomenon ought to be investigated to ensure models remain helpful and trustworthy.
@@
Contributions. To address this need, this work theoretically derives and experimentally evaluates the long-term evolutionary behavior of generative models trained on each other’s data. Specifically, we
@@
Both our theoretical analysis and experiments show that when training with a mixture of real and synthetic data, the implicit interaction between heterogeneous models and datasets can have both positive and negative impacts. At a high level, well known concepts in statistical learning theory anticipate this: recursive training on the same data is bad (e.g., overfitting and model “collapse”) but training on novel data, even if synthetic, can boost performance (e.g., transfer learning). Our experimental results provide concrete evidence that these phenomena can occur simultaneously.
@@ 2 Related Work
Model collapse is a recently observed phenomenon in large-scale generative text and image models. It referred—in it’s earliest form—to the phenomenon of models performing much worse after generations of training on their own generated outputs (Shumailov et al., 2024; Peterson, 2025; Wang et al., 2024; Alemohammad et al., 2024; Feng et al., 2024; Hataya et al., 2023; Dohmatob et al., 2025; Martínez et al., 2023). Theoretical and empirical results from these works show that models, if trained on generated outputs from their prior versions, slowly degrade in performance as generations progress. One way this manifests is in models forgetting the tails of their original (real) training data, since generated content tends not to contain rare content from the original training data. Training repeatedly on truncated, synthetic data leads the model to forget the richness of its original distribution, resulting in degraded behavior (at best) and total failure (at worst).
Mitigating model collapse. Despite the dire predictions of these papers, subsequent work has proposed a simple mitigation strategy: instead of discarding all prior (human-generated) training data, retain some fraction of this while augmenting it with generated data. Numerous works have observed that this choice to augment instead of discard the original training dataset results in a bounded error in future models, avoiding collapse (Kazdan et al., 2025; Gerstgrasser et al., 2024; Marchi et al., 2024). Although most of these results were discovered on small models, recent work claims that the observed bound in error $\pi^{2}/6$ exists for all models (Dey & Donoho, 2024). Further work (Schaeffer et al., 2025) summarizes current research on collapse.
Transfer learning and other model interactions. Significant prior work has studied transfer learning, in which information learned by one model is passed to another, often by reusing the trained weights of a “teacher” model to initialize a “student” model (Zhuang et al., 2020). Some prior work has further considered the use of synthetic data in transfer learning (Tian & Shen, 2025; Kim et al., 2022; Brinner et al., 2025). Our work is distinct from transfer learning due to its focus on unintentional data-mediated interactions between models. Furthermore, limited work has examined long-term effects of models training on each other’s data.  Zhang et al. (2024) consider the setting where a generative model is trained on data generated by other models, but does not consider long-term effects of such interactions among multiple models.  Jain & Krishnamurthy (2025) study interacting Large Language Model agents through the lens of Bayesian social learning and microeconomics, but do not focus specifically on data-mediated interactions between models.
@@ 3 How Today’s Large-Scale Generative AI Models Are Trained
Step 1: Pretraining. Following well-established scaling laws linking model performance and dataset size (Kaplan et al., 2020), today’s large-scale generative AI models are trained on massive datasets often scraped from the internet. Early versions of GPT, Llama, and PaLM all report being trained on scraped datasets like Common Crawl, ArXiv, Github, Wikipedia, and/or Stack Exchange (Touvron et al., 2023a; Chowdhery et al., 2023; Brown et al., 2020)—see Table 1 for an overview.
Table 1: Examples of training data listed for prominent language models. = use explicitly stated; = significant overlap expected (e.g. GLaM (Du et al., 2022) and PaLM (Chowdhery et al., 2023) are trained on Microsoft’s internal re-creation of WebText (Brown et al., 2020)). We only include models for which training data sources are explicitly stated. See Table 5 in Appendix for information on other prominent models.
@@
Chinchilla (Hoffmann et al., 2022)
GLaM (Du et al., 2022)
GPT (Radford et al., 2018)
GPT-2 (Radford et al., 2019)
GPT-3 (Brown et al., 2020)
LaMDA (Thoppilan et al., 2022)
Llama 1 (Touvron et al., 2023a)
PaLM (Chowdhery et al., 2023)
Phi 2 (Abdin et al., 2023)
@@
Another striking fact emerges from the categorization of training data in Table 1: large-scale model training datasets overlap. For example, GPT, Jamba, Llama, PaLM, and Phi are all trained on subsets of CommonCrawl (Crawl, 2025), while GPT, Llama, and PaLM are all trained on Wikipedia and Books datasets. Several other models have other points of training data overlap. This suggests that:
@@
(Touvron et al., 2023a):
ArXiv, Books, Common Crawl,
@@
(Dubey et al., 2024):
"A variety of data sources containing
@@
GPT
(Radford et al., 2018):
BooksCorpus
(Radford et al., 2019):
WebText
(Brown et al., 2020):
CommonCrawl, WebText2,
@@
(Li et al., 2023):
The Stack, Stack Overflow,
@@
(Abdin et al., 2023):
The Stack, Stack Overflow,
@@
(Abdin et al., 2024):
"publicly available web data…
@@
Step 2: Fine-tuning. Variously called fine-tuning or alignment, this phase leverages proprietary methods or data to tweak model behaviors in ways model providers believe are helpful. For example, the LLama fine-tuning phase (Dubey et al., 2024) involves many rounds of reinforcement learning with human feedback (RLHF) to stamp out model negative behaviors, while Phi (Abdin et al., 2024) was fine-tuned on proprietary synthetic data to patch “gaps” in its mathematical reasoning abilities.
Step 3: Model updates. A key assumption of prior literature on model collapse is that models are updated. Updating generally involves training a new model using similar architecture initialized with weights from the prior step, techniques, and datasets to prior versions of the model. This matches reality—the literature documents numerous “families” of models that directly descend from one another (c.f. LLama 1, 2, and 3 (Touvron et al., 2023a; b; Dubey et al., 2024); GPT 1, 2, 3, 4 (Radford et al., 2018; Brown et al., 2020; Radford et al., 2019; Achiam et al., 2023); Phi 1.5, 2, 3 (Li et al., 2023; Abdin et al., 2023; 2024)). Prior works on model collapse have proposed several ways training data from prior model generations could be re-used (or not) during model updates.
The replace scenario, first proposed by Shumailov et al. (2024) assumes model trainers completely replace their training data at each update, using only generated data output by the prior version of the model. Schaeffer et al. (2025) point out that this edge case is interesting but impractical. The accumulate scenario (Gerstgrasser et al., 2024) assumes model trainers augment their original data with additional data that may contain AI-generated content at each update set. Finally, the accumulate and subsample scenario (Kazdan et al., 2025) echoes the accumulate scenario but subsamples a fixed-sized subset from the original and accumulated data at each update step (rather than letting dataset size grow linearly). This scenario acknowledges the real-world compute limitations.
We believe that the accumulate-and-subsample paradigm best reflects reality, so we will leverage it in our work. We support this opinion with evidence from three well-documented model families: Llama, GPT, and Phi. Table 2 records the training data used in publicly disclosed generations of these models. As the table shows, trainers re-use some prior training data for model updates, supplementing this with additional web content. Whitepapers for models published after 2023 do not contain specific training data information but suggest collection of additional online data at each update.
@@
However, most prior work on model collapse still overlooks a fundamental reality in model updates: future models trained on internet-sourced content will be trained on outputs from other generative AI models, not merely their own. Already, the internet is filled with generated content from various models (Sun et al., 2025). As model trainers collect new data to facilitate model updates, internet-sourced data will inevitably contain content from other generative models.
@@ 4 Formalizing An Iterative, Interactive Model Training Pipeline
Section 3 provides empirical evidence for two realities overlooked by prior studies: internet-scraped training datasets used for initial training may have substantial overlap, and models may be updated using each others’ generated outputs. To study the effect of these two factors on model evolution, we propose a general workflow in which multiple entities regularly update their models using a mix of private, public, and generated data. Based on §3, we consider three types of training/update data:
@@
  • $D_{t}=\{D_{t1},D_{t2},\ldots D_{tk}\}$: Public data used for updates at time $t$ by multiple entities (synthetic data). $D_{tk}$ is data generated by the $k^{th}$ entity based on model $\theta_{t-1,k}$.
@@
  • $\alpha$, $0\leq\alpha\leq 1$ is the fraction of new data introduced at generation $t$, relative to the amount of initial data reused (following the “accumulate and subsample” paradigm of Kazdan et al. (2025)).
@@
    Each entity $k=1,\dots,K$ updates it model parameters via $\hat{\theta}_{t,k}=\Phi_{t,k}(\hat{\theta}_{t-1,k},D_{*},\tilde{D}_{k},D_{t})$. Here, $\Phi_{0,k}$ is a training algorithm that depends on the previous model parameter $\hat{\theta}_{k,t-1}$ as well as the data. As before, training may employ subsampling, weighting, and randomization.
@@ 5 Theory
We theoretically analyze the behavior of the interactive workflow. Similar to prior work (Gerstgrasser et al., 2024; Kazdan et al., 2025; Dey & Donoho, 2024; Dohmatob et al., 2025; Barzilai & Shamir, 2025), we focus on the linear regression models where each data point consists of a feature-response pair $(x,y)\in\mathbb{R}^{d}\times\mathbb{R}$. By the universality results of Dey & Donoho (2024), the analysis of this settings also applies to generalized linear models satisfying appropriate asymptotic normality assumptions.
@@
We note that our framework could easily be extended to accommodate new, human-generated data at each time step, but we omit this in our formulation for analytic simplicity. Throughout our analysis we assume that all features are deterministic. We represent dataset $\tilde{D}_{k}$ with $\tilde{n}_{k}\times d$ matrix $\tilde{X}_{k}=[\tilde{x}_{k1},\dots,\tilde{x}_{k\tilde{n}_{k}}]^{\top}$ and $\tilde{n}_{k}\times 1$ vector $\tilde{y}_{k}=[\tilde{y}_{k1},\dots,y_{k\tilde{n}_{k}}]^{\top}$, and use the same convention for the public data $(X_{*},y_{*})$ and the generated data $(X_{tk},y_{tk})$. Data across different entities are then combined into “lifted” representations, which are denoted using boldface:
@@
Classical results in matrix analysis (Higham, 2002) imply that if the spectral radius of $\bm{Q}$ is strictly less than one, then these these matrices converge to well-defined limits $\bm{M}$ and $\bm{C}$ satisfying
@@
The following result provides a sufficient condition for convergence in terms of the triple $(\tilde{\bm{S}},S_{*},\bm{S})$. In particular, if $\bm{S}$ is proportional to $\tilde{\bm{S}}$ then the condition is satisfied for all $0\leq\alpha<1$ and $0\leq\beta\leq 1$. Note that the boundary case $\alpha=1$ corresponds to the recursive training setting of Shumailov et al. (2024) where the variance increases linearly across generations, and thus convergence does not occur.
@@
Suppose that $\bm{S}\propto\lambda\tilde{\bm{S}}+(1-\lambda){\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}(\mathrm{I}_{K}\otimes S_{*})}$ for some $0<\lambda\leq 1$. Then, the spectral radius of $\bm{Q}$ is strictly less than one for all $0\leq\alpha<1$ and $0<\beta\leq\lambda$.
@@
We can use this to compute the optimal $\alpha$, $\beta$ values for the interactive training setting. Figure 2 compares the asymptotic MSE for a training workflow with given $\alpha$, $\beta$ (from Theorem 3) with the MSE of an idealized setting where each entity has access to the entire collection of real data (both private and public). Each curve represents the relative efficiency, i.e., the ratio of optimal MSE to entity-specific workflow MSE, with values close to one indicating near optimality. These results demonstrate that a setting with $\beta=0.5$ and $\alpha=0.5$ achieves the best global performance for all models. When $\beta$ is much larger (0.9), relatively small $\alpha$ values also improve model performance.
@@
Figure 2: Predicted relative efficiency across $\alpha,\beta$ values for a $K=4$ model system (dimension $15$ and rank 55). Curves show ratio of MSE of the minimum variance unbiased estimator to asymptotic MSE for a given $\alpha,\beta$ setting obtained from Theorem 3 (11 is optimal). Setting $\alpha=\beta=0.5$ produces best results across models.
6 Experimental Evaluation
To understand how our theoretical predictions bear out in practice, we run experiments on text-generation models. In each, we train $K$ interacting models (per our framework in Figure 1) for several generations and evaluate how their performance changes on their own and other models’ tasks. Here, we present the setup and results for a $K=2$ model system. Additional results are in the Appendix.